Administer 1000 ml of 0.45 NS over 8 hrs using 10 gtts / ml drip set. What is the gravity infusion rate?
Answer #1

Answer :

Use the intravenous flow rate formula:

Total volume ×Drop factor ÷Times in minute =Drops per minute

HERE:

Total volume =1000ml

Drop factor=10gtt

Time per minute =60×8=480minutes

The gravity infusion rate is

1000×10gtt÷480=20.83gtt/minute